# Nonlocal Josephson diode effect in minimal Kitaev chains

### v1

@@ Abstract
Critical currents and nonlocal Josephson diode effect.—To further explore the nonlocal nonreciprocal Josephson transport, we look at the critical currents across the left junction $I_{\rm cL}^{\pm}(\phi_{\rm R})={\rm max}_{\phi_{\rm L}}[\pm I_{\rm L}(\phi_{\rm L},\phi_{\rm R})]$, where $I_{\rm L}(\phi_{\rm L},\phi_{\rm R})$ is the current flowing across the left junction discussed in the previous section. Figs. 4(a-c) show the critical currents $I_{\rm cL}^{\pm}(\phi_{\rm R})$ as a function of $\phi_{\rm R}$ when ECT in the middle chain is weaker, equal, and stronger than CAR in the same chain. At $t_{\rm M}=\Delta_{\rm M}$, when CAR and ECT in the middle chain are distinct, the positive and negative critical currents are equal ( $I_{\rm cL}^{+}(\phi_{\rm R})=I_{\rm cL}^{-}(\phi_{\rm R})$) and develop a minimum at $\phi_{\rm R}=\pi$ [Fig. 4(b)], leading to the absence of the nonlocal Josephson diode effect. Interestingly, for $t_{\rm M}<\Delta_{\rm M}$, the minima of $I_{\rm cL}^{+}(\phi_{\rm R})$ and $I_{\rm cL}^{-}(\phi_{\rm R})$ are shifted to values below and above $\phi_{\rm R}=\pi$, respectively, giving rise to nonreciprocal phase-biased Josephson currents ($I_{\rm cL}^{+}(\phi_{\rm R})\neq I_{\rm cL}^{-}(\phi_{\rm R})$) flowing across the left junction and fully controlled by $\phi_{\rm R}\neq n\pi$. For $t_{\rm M}>\Delta_{\rm M}$, the shifts in the critical currents is the opposite as for $t_{\rm M}<\Delta_{\rm M}$ but $I_{\rm cL}^{+}(\phi_{\rm R})\neq I_{\rm cL}^{-}(\phi_{\rm R})$ remains. Having $I_{\rm cL}^{+}(\phi_{\rm R})\neq I_{\rm cL}^{-}(\phi_{\rm R})$ defines the nonlocal Josephson diode effect, which requires an imbalance between CAR and ECT in the middle chain and also $\phi_{\rm R}\neq n\pi$. These conditions are consistent with the asymmetric Andreev spectrum [Fig. 2] and with the breaking of local TR and charge-conjugation symmetries discussed below Eq. (2), implying that they are crucial for the nonlocal Josephson diode effect in minimal Kitaev chains.
@@
To characterize the nonlocal diode effect, we analyze its efficiency by studying the quality factor $\eta_{\rm L}(\phi_{\rm R})=[I_{\rm cL}^{+}(\phi_{\rm R})-I_{\rm cL}^{-}(\phi_{\rm R})]/[I_{\rm cL}^{+}(\phi_{\rm R})+I_{\rm cL}^{-}(\phi_{\rm R})]$, where $I_{\rm cL}^{\pm}(\phi_{\rm R})$ are the critical currents across the left junction analyzed in the previous paragraph. We show $\eta_{\rm L}(\phi_{\rm R})$ in Fig.  4(d,e) as a function of $\phi_{\rm R}$ and $t_{\rm M}$, while in Fig.  4(f,g) we inspect the stability of the efficiencies when varying the onsite energies and tunneling across the junction since these conditions are very likely unavoidable in a realistic scenario. In S3 and Fig. 7 of the End Matter, we discuss the robustness of $\eta_{\rm L}(\phi_{\rm R})$ against variations of the onsite energies $\varepsilon_{\alpha}$. In Fig.  4(d,e), we observe that a nonzero quality factor appears when $t_{\rm M}\neq\Delta_{\rm M}$ and $\phi_{\rm R}\neq n\pi$, with $n\in\mathbb{Z}$, and it is enhanced just below and above $\phi_{\rm R}=\pi$ achieving values above $0.5$. While $t_{\rm M}/\Delta_{\rm M}<1$ promotes a quality factor having a sine-line profile with $\phi_{\rm R}$ with a fast sign change across $\phi_{\rm R}=\pi$, the quality factors change sign for $t_{\rm M}/\Delta_{\rm M}>1$, see Fig.  4(d,e). This implies that the diode’s polarity and large efficiencies can be fully controlled nonlocally by $\phi_{\rm R}$ and is thus expected to be more tunable than regular two-terminal Josephson diodes [85, 86, 87, 88, 89, 90, 91, 69, 92, 93, 94, 95, 96, 97, 98, 68, 99, 100, 101]. The large efficiencies of the nonlocal Josephson diode effect are robust against variations of the onsite energies of the middle chain, as demonstrated in Figs. 4(f). Strong variations of the onsite energies in the left and right chains, however, have a detrimental effect although $\eta_{\rm L}(\phi_{\rm R})$ remains robust against weak detuning, see Fig. 7. Moreover, having large tunneling amplitudes favours the large efficiency values of the nonlocal Josephson diode [Fig. 4(g)]. We have also verified that slight deviations from the equal CAR and ECT regime in the left /right chains weakly affects the nonlocal Josephson diode effect.
@@
In conclusion, we have demonstrated the nonlocal Josephson effect and the realization of a nonlocal Josephson diode in a system formed by three laterally coupled minimal Kitaev chains. In particular, we found that when the middle chain exhibits an imbalance between crossed Andreev reflections and electron cotunneling, the phase-dependent Andreev spectrum is asymmetric and can be controlled by the phase difference across the right junction. We have then shown that this asymmetric Andreev spectrum, containing hybridized Andreev bound states from both junctions, produces a supercurrent across the left junction that is controllable by the phase difference across the right junction, hence defining the nonlocal Josephson effect. Notably, we have discovered that these nonlocal Josephson effects enable distinct positive and negative critical currents, giving rise to a nonlocal Josephson diode effect whose polarity is highly tunable and can achieve efficiencies above $50\%$. Given the recent advances on nonlocal Josephson effect [36, 37, 38, 39, 40, 41, 42] and minimal Kitaev chains [59, 60, 61, 62], our findings are very likely within experimental reach.
@@
J. C. acknowledges financial support from the Swedish Research Council (Vetenskapsrådet Grant No. 2021-04121). M. S. was supported by JST CREST Grant No. JPMJCR19T2 and JSPS KAKENHI Grant Nos. JP24K00569 and JP25H01250.

### v2

@@ VI Critical currents and nonlocal Josephson diode effect
To further explore the nonlocal nonreciprocal Josephson transport, we look at the critical currents across the left junction $I_{\rm cL}^{\pm}(\phi_{\rm R})={\rm max}_{\phi_{\rm L}}[\pm I_{\rm L}(\phi_{\rm L},\phi_{\rm R})]$, where $I_{\rm L}(\phi_{\rm L},\phi_{\rm R})$ is the current flowing across the left junction discussed in the previous section. Figs. 4(a-c) show the critical currents $I_{\rm cL}^{\pm}(\phi_{\rm R})$ as a function of $\phi_{\rm R}$ when ECT in the middle chain is weaker, equal, and stronger than CAR in the same chain. At $t_{\rm M}=\Delta_{\rm M}$, when CAR and ECT in the middle chain are equal, the positive and negative critical currents are equal ($I_{\rm cL}^{+}(\phi_{\rm R})=I_{\rm cL}^{-}(\phi_{\rm R})$) and develop a minimum at $\phi_{\rm R}=\pi$ [Fig. 4(b)], leading to the absence of the nonlocal Josephson diode effect. Interestingly, for $t_{\rm M}<\Delta_{\rm M}$, the minima of $I_{\rm cL}^{+}(\phi_{\rm R})$ and $I_{\rm cL}^{-}(\phi_{\rm R})$ are shifted to values below and above $\phi_{\rm R}=\pi$, respectively, giving rise to nonreciprocal phase-biased Josephson currents ($I_{\rm cL}^{+}(\phi_{\rm R})\neq I_{\rm cL}^{-}(\phi_{\rm R})$) flowing across the left junction and fully controlled by $\phi_{\rm R}\neq n\pi$. For $t_{\rm M}>\Delta_{\rm M}$, the shifts in the critical currents is the opposite as for $t_{\rm M}<\Delta_{\rm M}$ but $I_{\rm cL}^{+}(\phi_{\rm R})\neq I_{\rm cL}^{-}(\phi_{\rm R})$ remains. Having $I_{\rm cL}^{+}(\phi_{\rm R})\neq I_{\rm cL}^{-}(\phi_{\rm R})$ defines the nonlocal Josephson diode effect, which requires an imbalance between CAR and ECT in the middle chain and also $\phi_{\rm R}\neq n\pi$. These conditions are consistent with the asymmetric Andreev spectrum [Fig. 2] and with the breaking of local TRS and local charge-conjugation symmetry discussed below Eq. (2), implying that they are crucial for the nonlocal Josephson diode effect in minimal Kitaev chains.
@@
To characterize the nonlocal diode effect, we analyze its efficiency by studying the quality factor $\eta_{\rm L}(\phi_{\rm R})=[I_{\rm cL}^{+}(\phi_{\rm R})-I_{\rm cL}^{-}(\phi_{\rm R})]/[I_{\rm cL}^{+}(\phi_{\rm R})+I_{\rm cL}^{-}(\phi_{\rm R})]$, where $I_{\rm cL}^{\pm}(\phi_{\rm R})$ are the critical currents across the left junction analyzed in the previous paragraph. We show $\eta_{\rm L}(\phi_{\rm R})$ in Fig.  4(d,e) as a function of $\phi_{\rm R}$ and $t_{\rm M}$, while in Fig.  4(f,g) we inspect the stability of the efficiencies when varying the onsite energies and tunneling across the junction since these conditions are very likely unavoidable in a realistic scenario. In Appendix C, we discuss the robustness of $\eta_{\rm L}(\phi_{\rm R})$ against variations of the onsite energies $\varepsilon_{\alpha}$, see Fig. 7. In Fig.  4(d,e), we observe that a nonzero quality factor appears when $t_{\rm M}\neq\Delta_{\rm M}$ and $\phi_{\rm R}\neq n\pi$, with $n\in\mathbb{Z}$, and it is enhanced just below and above $\phi_{\rm R}=\pi$ achieving values above $0.5$. While $t_{\rm M}/\Delta_{\rm M}<1$ promotes a quality factor having a sine-line profile with $\phi_{\rm R}$ with a fast sign change across $\phi_{\rm R}=\pi$, the quality factors change sign for $t_{\rm M}/\Delta_{\rm M}>1$, see Fig.  4(d,e). This implies that the diode’s polarity and large efficiencies can be fully controlled nonlocally by $\phi_{\rm R}$ and is thus expected to be more tunable than regular two-terminal Josephson diodes [92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 75, 102, 103, 104, 74, 105, 106, 107, 108]. By direct inspection of Fig.  4(e), one concludes that, for $t_{\rm M}<\Delta_{\rm M}$, the largest efficiencies are achieved for $t_{\rm M}\in(0.3-0.7)\Delta_{\rm M}$ and $0.3\pi<\phi_{\rm R}<\pi$ ($\pi<\phi_{\rm R}<0.7\pi$). The large efficiencies of the nonlocal Josephson diode effect are robust against variations of the onsite energies of the middle chain, as demonstrated in Figs. 4(f). Strong variations of the onsite energies in the left and right chains, however, have a detrimental effect, while $\eta_{\rm L}(\phi_{\rm R})$ remains robust against weak detuning, see Fig. 7. Moreover, having large tunneling amplitudes favours the large efficiency values of the nonlocal Josephson diode [Fig. 4(g)]. We have also verified that slight deviations from the equal CAR and ECT regime in the left /right chains weakly affects the nonlocal Josephson diode effect.
VII Conclusions
@@
In conclusion, we have demonstrated the nonlocal Josephson effect and the realization of a nonlocal Josephson diode in a system formed by three laterally coupled minimal Kitaev chains. In particular, we found that when the middle chain exhibits an imbalance between crossed Andreev reflections and electron cotunneling, the phase-dependent Andreev spectrum is asymmetric and can be controlled by the phase difference across the right junction. We have then shown that this asymmetric Andreev spectrum, containing hybridized Andreev bound states from both junctions, produces a supercurrent across the left junction that is controllable by the phase difference across the right junction, hence defining the nonlocal Josephson effect. Notably, we have discovered that these nonlocal Josephson effects enable distinct positive and negative critical currents, giving rise to a nonlocal Josephson diode effect whose polarity is highly tunable and can achieve efficiencies above $50\%$. Given the recent advances on nonlocal Josephson effect [38, 39, 40, 41, 42, 43, 44] and minimal Kitaev chains [62, 63, 64, 65], our findings are very likely within experimental reach. In this regard, we point out that Ref. [62] has already experimentally demonstrated the control of the ratio between crossed Andreev reflections and electron cotunneling, which, applied to the middle chain in our setup, shows that the realization of the nonlocal Josephson diode effect with a highly tunable polarity in minimal Kitaev chains is a feasible idea. While our work makes us wonder about emergent effects in minimal Kitaev chains, an immediate question also involves the application of the nonlocal Josephson diode concept in longer Kitaev chains, which will be the focus of our future research.
Acknowledgements.
@@
J. C. acknowledges financial support from the Swedish Research Council (Vetenskapsrådet Grant No. 2021-04121) and from the Olle Engkvist Foundation (Grant No. 243-1026). M. S. was supported by JST CREST Grant No. JPMJCR19T2 and JSPS KAKENHI Grant Nos. JP24K00569 and JP25H01250. The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through Grant Agreement No. 2022-06725.
Appendix A Andreev bound states when ECT is stronger than CAR in the middle chain
In Fig. 2 we looked at the Andreev spectrum when ECT is equal to or weaker than CAR in the middle chain. In this part, Fig. 5 shows the Andreev spectrum as a function of $\phi_{\rm L}$ for distinct $\phi_{\rm R}$ when ECT is stronger than CAR, namely, $t_{\rm M}>\Delta_{\rm M}$. The Andreev spectrum is symmetric when $\phi_{\rm R}=0$ [Fig. 5(a)], developing a phase dependence similar to the case with $t_{\rm M}<\Delta_{\rm M}$ in Fig. 2(a). Because we chose $\Delta_{\rm L,R}=t_{\rm L,R}$ and $\varepsilon_{\alpha}$, the spectrum contains two dispersionless levels at zero energy, which represent the two poor man’s Majorana modes located at the very outer QDs. Two extra Andreev bound states reach zero energy at $\phi=\pi$ marked by a vertical dashed red line, and those are located in the middle of the junction. However, deviations from these parameters do not change the symmetric nature of the spectrum in Fig. 5(a) unless $\phi_{\rm R}\neq n\pi$. This is demonstrated in Fig. 5(b-d), where a $\phi_{\rm R}\neq n\pi$ makes the entire Andreev spectrum asymmetric with respect to $\phi_{\rm L}=\pi$, see vertical dashed gray line. The zero-energy crossing seen at $\phi_{\rm L}=0$ in Fig. 5(a) shifts to $\phi_{\rm L}>\pi$ when $\phi_{\rm R}<\pi$ [Fig. 5(b)], while to $\phi_{\rm L}<\pi$ for $\phi_{\rm R}>\pi$ [Fig. 5(d)], see vertical dashed red lines. At $\phi_{\rm R}=\pi$, the Andreev specturm is symmetric, the pair of dispersionless levels remain, but the first excited levels now reach zero energy at $\phi_{\rm L}=2n\pi$ while leaving a large zero energy splitting at $\phi_{\rm L}=n\pi$, Fig. 5(c). In conclusion, the asymmetry in the Andreev spectrum is controlled by $\phi_{\rm R}$ and $\Delta_{\rm M}\neq t_{\rm M}$, in line with the breaking local time-reversal and local charge-conjugation symmetries discussed in the main text.
Figure 5: Energy spectrum as a function of $\phi_{\rm L}$ at distinct values of $\phi_{\rm R}$ for $t_{\rm M}=1.8\Delta_{\rm M}$. The vertical dotted line marks $\phi_{\rm L}=\pi$, while the red vertical dashed line marks the value of $\phi_{\rm L}$ at which the dispersing lowest ABSs reach zero energy. Parameters: $\varepsilon_{\alpha}=0$, $t_{\rm L,R}=\Delta_{\alpha}\equiv\Delta=1$, $\tau_{\rm L}=1$, $\tau_{\rm R}=0.9$.
Appendix B Free energy of the Josephson setup
The asymmetric Andreev spectrum showed in the previous section has a direct effect on the free energy, which we discuss in this part. We define a normalized free energy as $\bar{F}=(F-F_{\rm min})/(F_{\rm max}-F_{\rm min})$, where $F_{\rm max}={\rm max}_{\phi_{\rm L}}[F(\phi_{\rm L},\phi_{\rm R})]$ and $F_{\rm min}={\rm min}_{\phi_{\rm L}}[F(\phi_{\rm L},\phi_{\rm R})]$, where the maximum and minimum are obtained over a period of $2\pi$. In Fig. 6, we show $\bar{F}$ as a function of $\phi_{\rm L}$ for distinct values of $\phi_{\rm R}$ when ECT in the middle chain $t_{\rm M}$ is equal, weaker, and stronger than CAR in the same chain $\Delta_{\rm M}$. As anticipated, for $t_{\rm M}=\Delta_{\rm M}$ in Fig. 6(a), the free energy $\bar{F}$ develops minima at multiples of $2n\pi$, with $n\in\mathbb{Z}$, irrespective of the values of $\phi_{\rm R}$. This signals a regular Josephson junction behavior. Interestingly, for $t_{\rm M}\neq\Delta_{\rm M}$ and $\phi_{\rm R}\neq 0$, the free energy exhibits minima away from $\phi_{\rm L}=2n\pi$, see Fig. 6(b,c). For instance, for $\phi_{\rm R}=\pi$ at $t_{\rm M}\neq\Delta_{\rm M}$, the free energy has two (double) minima away from $\phi_{\rm L}=0,\pi$, signaling the emergence of a $\phi$-Josephson junction. We note that $\phi$-Josephson junctions were initially predicted in Josephson junctions with alternating critical current density [88], in periodic arrays of 0- and $\pi$-Josephson junctions [89], and more recently in altermagnetic Josephson junctions [90, 91, 25]. Furthermore, away from $\phi_{\rm R}=\pi$ at $t_{\rm M}\neq\Delta_{\rm M}$, the free energy is able to exhibit one minimum either close to $\phi_{\rm L}=0$ or $\phi_{\rm L}=2\pi$: on one hand, when $\pi<\phi_{\rm R}<2\pi$, the minimum is close to $\phi_{\rm L}=0$ for $t_{\rm M}<\Delta_{\rm M}$ but close to $\phi_{\rm L}=2n\pi$ for $t_{\rm M}>\Delta_{\rm M}$. On the other hand, when $0<\phi_{\rm R}<\pi$, the minimum is close to $\phi_{\rm L}=2\pi$ for $t_{\rm M}<\Delta_{\rm M}$ but close to $\phi_{\rm L}=0$ for $t_{\rm M}>\Delta_{\rm M}$. The asymmetric Andreev spectrum and nonreciprocal (diode) Josephson transport discussed in the main text is therefore tied to the free energy having a single minimum away from $\phi_{\rm L}=2n\pi$.
Figure 6: Free energy $\bar{F}$ as a function of $\phi_{\rm L}$ at distinct values of $\phi_{\rm R}$ for $t_{\rm M}=\Delta_{\rm M}$, $t_{\rm M}=0.5\Delta_{\rm M}$, and $t_{\rm M}=1.8\Delta_{\rm M}$. Here, $\bar{F}$ is defined as $\bar{F}=(F-F_{\rm min})/(F_{\rm max}-F_{\rm min})$, where $F$ is the free energy and $F_{\rm max(min)}$ is the maximum (minimum) free energy value over a period of $\phi_{\rm L}$. Parameters: $\varepsilon_{\alpha}=0$, $t_{\rm L,R}=\Delta_{\alpha}\equiv\Delta=1$, $\tau_{\rm L}=1$, $\tau_{\rm R}=0.9$.
Appendix C Efficiency of the nonlocal Josephson diode at finite onsite energies
To further support the robustness of the nonlocal Josephson diode effect presented in Fig. 4, here we explore the sensitivity of the quality factor $\eta_{\rm L}$ under variations of the onsite energies $\varepsilon_{\alpha}$. This is presented in Fig. 7(a,b), where we show $\eta_{\rm L}$ as a function of $\phi_{\rm R}$ when the onsite energies of the first QD of the left and right chains take finite values, while in Fig. 7(b,c) we consider the case when both onsite energies of the left and right chains are nonzero and equal. The quality factors exhibit a slightly greater robustness against variations of a single onsite energy [Fig. 7(a,b)], achieving values close to $20\%$ for detuning values of $\epsilon_{1}/\Delta=0.5$. When both QD energies of the outer chains equally large, the quality factors degrade faster than when a single QD has a nonzero energy; see Fig. 7(c,d).
While these QD energies seem to be detrimental, it is worth noting that they also bring additional functionality for the nonlocal Josephson diode. To see this, we first notice that the polarity of the diode, reflected in the sign of $\eta_{\rm L}$, is highly controllable by the phase across the junction between the middle and right chains $\phi_{\rm R}$. The diode’s polarity is also determined by strength of ETC and CAR, as observed in Fig. 7. Notably, the nonzero values of the onsite energies in the outer chains have the capability to change the diode’s polarity around $\phi_{\rm R}=\pi$. Depending on whether a single $\epsilon_{1}$ or both $\epsilon_{1,2}$ energies are varied, the change in the diode’s polarity can be fast or slow, respectively. Since the QD energies are expected to be tunable by means of gates [62], $\epsilon_{1,2}$ add important control of the nonlocal Josephson diode effect discussed here.
Figure 7: (a,b) Diode’s efficiency as a function of $\phi_{\rm R}$ for distinct values of the onsite energy $\varepsilon_{\rm L_{1},R_{1}}=\epsilon_{1}$ when $t_{\rm M}=0.5\Delta_{\rm M}$ (a) and $t_{\rm M}=1.8\Delta_{\rm M}$ (a), in both cases at $\varepsilon_{\rm L_{2}(R_{2})}=0$. (c,d) The same as in (a,b) but both onsite energies $\varepsilon_{\rm L_{1,2}}=\varepsilon_{\rm R_{1,2}}=\epsilon_{1,2}$. Parameters: $\varepsilon_{\rm M}=0$, $t_{\rm L,R}=\Delta_{\alpha}\equiv\Delta=1$, $\tau_{\rm L}=1$, $\tau_{\rm R}=0.9$.